OpenStudy (christos):

\[\sqrt{{{a}^{2}}-{{x}^{2}}}\text{ Help me solve this thing please.}\]

OpenStudy (christos):

ops, its supposed to be an integral.

OpenStudy (christos):

\[\int{{}}(\sqrt{{{a}^{2}}-{{x}^{2}})}\] there

OpenStudy (anonymous):

so here we have a function inside of a function like this: f(g(x)) after changing the radical into an exponent, you will want to use the chain rule ^_^

OpenStudy (christos):

I am afraid I am too inexperienced to follow through your logic :S

OpenStudy (anonymous):

no problem! can yo do the first step, change the radical into an exponent?

OpenStudy (christos):

\[\int{(}{{A}^{2}}-{{x}^{2}}{{)}^{(1/2)}}\]

OpenStudy (christos):

ok ??

OpenStudy (loser66):

let x = asin u \(\rightarrow dx = acos u du\) x^2 = a^2 sin^2 u a^2 - x^2 = a^2 - a^2sin^2 u get this part?

OpenStudy (loser66):

no??

OpenStudy (loser66):

still have some steps to get the answer, but if you don't get this part I am better stop here. Good luck

OpenStudy (christos):

can you help me with this part here ? a^2 - x^2 = a^2 - a^2sin^2 u

OpenStudy (christos):

nvm I got everything up until now @Loser66

OpenStudy (christos):

\[\int{{}}\sqrt{({{a}^{2}}-{{a}^{2}}si{{n}^{2}}theta)}\]

OpenStudy (loser66):

so?

OpenStudy (loser66):

you miss cos \(\theta\)d\(\theta\)

OpenStudy (loser66):

I use u and you use \(\theta\), hopefully you can see they are just notation.

OpenStudy (christos):

and then ?

OpenStudy (loser66):